OpenStudy (anonymous):

If the perimeter of an isosceles trapezoid is 59.9 meters, and the bases measure 23.2 meters and 17.1 meters, find the length of the legs of the trapezoid. A. 18.35 meters B. 9.8 meters C. 4.9 meters D. 19.6 meters

OpenStudy (anonymous):

2x + 23.2 + 17.1 = 59.9 solve for x = ?

OpenStudy (anonymous):

lol I don't know. how do I solve it?

OpenStudy (anonymous):

Tanjung has you covered :P

OpenStudy (anonymous):

2x + 23.2 + 17.1 = 59.9 2x +40.3 = 59.9 from both sides, subtract by 40.3 : 2x +40.3 - 40.3 = 59.9 - 40.3 simplify this, then divided by 2, get up your answer ...

OpenStudy (anonymous):

4.9?

OpenStudy (anonymous):

2x +40.3 - 40.3 = 59.9 - 40.3 2x = 19.6 2x/2 = 19.6/2 x = 9.8

OpenStudy (anonymous):

Thanks i understand now~!